Given a family with three children, find the probability of the event.
given that the birth order alternates between girls and boys. What is the probability that the youngest is a boy, given that the birth order alternates between girls and boys?

Answer by greenestamps(13203)



There are 2*2*2 = 8 possible boy/girl sequences.

Of those 8, only two are sequences in which the birth order alternates between boys and girls: BGB and GBG.

Of those two, only one has a boy as the youngest.

ANSWER: The probability that the youngest is a boy, given that the birth order alternates between boys and girls, is 1/2.

More formally, using the definition of conditional probability, theprobability that the youngest is a boy, given that the birth order alternates between boys and girls, is:

P((birth order alternates between boys and girls)AND(youngest is a boy)) (1/8) = 1
-------------------------------------------------------------------------- = ----- = ---
P(birth order alternates between boys and girls) (2/8) = 2